# Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model

### v2

@@ Abstract
While we have used the localization indicator $P$ (Supplementary Eq. (S59)) in the above numerical calculations, there are other possible definitions of the localization indicators. Specifically, we can consider the following definitions that distinguish the localization and anti-localization of zero modes $\Psi_{j}(x)$ in a global sense: (1) $P_{1}=w_{\rm edge}/w_{\rm ave}$ with $w_{\rm ave}=\sum_{x,i}|\Psi_{i}(x)|^{2}/L$ being the averaged amplitude, (2) $P_{2}=w_{\rm edge}/w(x)$ with $w(x)=\sum_{i}|\Psi_{i}(x)|^{2}$ being the amplitude and $x$ denoting a bulk site that we determine in hand, (3) $P_{3}=w_{\rm edge}/w_{\rm max}$ with $w_{\rm max}=\max_{x}\left(\sum_{i}|\Psi_{i}(x)|^{2}\right)$ being the maximum amplitude. Unfortunately, these localization indicators can be inconsistent with the nonlinear winding number at $a<2$, where the bulk-edge correspondence still holds true in the nonlinear SSH model. Supplementary Figure S4 shows the localization indicators of the nonlinear eigenvectors with the minimum absolute values of nonlinear eigenvalues. At $a=1.5$, all the indicators correspond to the localization and anti-localization of zero modes predicted from the nonlinear winding number. In contrast, at $a=1.9$, the indicators do not correspond to the nonlinear winding number. In this parameter regime, the bulk-edge correspondence is confirmed in semi-infinite systems (Fig. 2 in the main text). Therefore, the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$ are unsuitable to discuss the bulk-edge correspondence in finite nonlinear systems.
@@
.6 Linear stability analysis of the fixed points in the dynamical system describing zero modes of the extended nonlinear SSH model.
@@
Since the dynamical system in Eq. (9) in the main text can have three qualitatively different fixed points and limit cycles at $|\psi_{\rm fix}|^{2}=0$, $(d-a-\alpha)/b$, and $(\alpha-a)/b$, we below calculate the linear stability of these fixed points. We note that the nonzero fixed points, $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$ and $(\alpha-a)/b$, correspond to the topological edge modes that appear in the nonlinearity-induced topological phase because they converge to nonzero amplitude in the limit of $x\rightarrow\infty$. Below we focus on the parameter region $a,d,\alpha>0$, while we can conduct a similar analysis in the cases of different choices of the signs of the parameters.
.6.1 Stability analysis of the fixed point at $\psi_{\rm fix}=0$.
@@
Thus, one can determine the linear stability of the fixed point by comparing $a$ and $\alpha$. In the case of $a<\alpha$, the fixed point is linearly stable and has a two-dimensional stable manifold. In this parameter region, the winding number becomes $\nu_{\rm NL}=2$, which corresponds to the dimension of the stable manifold. In contrast, when $a$ is larger than $\alpha$, the fixed point is fully unstable. Then, the winding number becomes $\nu_{\rm NL}=0$, which is also equal to the dimension of the stable manifold.
@@
and thus confirm $|\lambda_{-}|>1$ and $|\lambda_{+}|<1$, which indicates that the stable manifold is one-dimension. In contrast, if we consider the parameter region $d<a+\alpha$, we obtain $d^{2}-4\alpha a<(d-2\alpha)^{2}$. Then, the sign of $2\alpha-d$ is determined by the sign of $\alpha-a$ as shown from the inequality
@@
Therefore, the dimension of the stable manifold is two in the case of $\alpha>a$ and zero in the case of $\alpha<a$. These results are consistent with the bulk-edge correspondence between the winding number and the dimension of the stable manifold.
.6.2 Stability analysis of the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$.
If we consider the parameter region $d<a+\alpha$, $\alpha<a$, and $b<0$, we obtain the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$. This fixed point corresponds to the topological edge modes that emerge by the nonlinearity-induced topological phase transition at $w=(d-a-\alpha)/b$. In this case, the eigenvalues of the matrix in Supplementary Eq. (S62) become
@@
We can show that these eigenvalues are real. Since the nonlinear winding number becomes $\nu_{\rm NL}=1$ at $w>(d-a-\alpha)/b$, we expect that the dimension of the stable manifold of the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$ is one. However, we can confirm the breakdown of such bulk-edge correspondence by the period-doubling bifurcation as in the original nonlinear SSH model in Eqs. (4) and (5) in the main text.
@@
Therefore, $|\lambda_{+}|$ becomes smaller than one and thus the nonlinear winding number corresponds to the dimension of the stable manifold only in the case of $(a+\alpha)/d>2$. At the critical parameter $(a+\alpha)/d=2$, the dynamical system exhibits a period-doubling bifurcation. Such bifurcation induces the breakdown of the bulk-edge correspondence as in the original nonlinear SSH model in Eqs. (4) and (5) in the main text.
.6.3 Stability analysis of the fixed point at $|\psi_{\rm fix}|^{2}=(\alpha-a)/b$.
If we consider the case of $2\alpha>d$ and $\alpha<a$, we obtain a limit cycle described as $\psi(x)=\sqrt{(\alpha-a)/b}\exp(i(\Omega x+\theta))$ whose frequency $\Omega$ satisfies $\Omega=-2\alpha\cos\theta$. This limit cycle corresponds to the topological edge mode that emerges by the nonlinearity-induced topological phase transition at $w=(\alpha-a)/b$, and the nonlinear winding number is changed into $\nu_{\rm NL}=2$ at the transition point. The eigenvalues of the linearized matrix in Supplementary Eq. (S62) become
@@
Thus, we obtain $|\lambda_{+}|<1$ only at $a-d/2<2\alpha$. If we consider the case of $a+d/2>2\alpha$, the dimension of the stable manifold is less than two, which indicates the breakdown of the bulk-edge correspondence. This breakdown is also induced by the bifurcation of the fixed point. Therefore, the breakdown of the bulk-edge correspondence by the chaos transition is independent of the hopping range in nonlinear models.
.7 Correspondence between the nonlinear winding number and the eigenvalues of a state-dependent transfer matrix.
We have also considered more general nonlinear systems whose zero modes are described by Eq. (11) in the main text. We here show that the eigenvalues of $T(f(\bm{\psi}_{1}(x)))$ in Eqs. (11) and (12) in the main text are related to the nonlinear winding number, which is the origin of the bulk-edge correspondence. Specifically, if we consider a fixed value of $f(\bm{\psi}(k))=w$ independently of the wavenumber $k$, the nonlinear winding number
@@
.8 Possible $\mathbb{Z}\times\mathbb{Z}$ classification in the nonlinear SSH model.
If we consider the case of $b\neq c$ in the nonlinear SSH model (Eqs. (4) and (5) in the main text), edge modes localized at the left and right boundaries can appear at different critical amplitudes. To show that, we consider both left and right semi-infinite systems, $x\leq-1$ and $x\geq 1$, respectively. While the spatial distribution of zero modes in the right semi-infinite system is described by
@@
(the same as Eq. (7) in the main text), that in the left semi-infinite system is described by
@@
and $\Psi_{A}(x)=0$. The difference between $b$ and $c$ leads to the different transition points in Supplementary Eqs. (S81) and (S82); the edge mode in the right semi-infinite system appears at $w=(d-a)/b$, while that in the left semi-infinite system appears at $w=(d-a)/c$. Therefore, the left edge mode and right edge mode have different critical amplitudes.
@@
whose transition points are consistent with the amplitude where the left and right edge modes appear. Such a $\mathbb{Z}\times\mathbb{Z}$ classification may be analogous to that in non-Hermitian systems, where the point-gap topology of one-dimensional systems is also classified by $\mathbb{Z}\times\mathbb{Z}$ topological invariants under the existence of the sublattice symmetry Kawabata2019 . However, the nonlinear SSH model considered here is a conservative system and preserves the energy in its time evolution. Therefore, nonlinearity in itself can induce non-Hermitian-like effects in the classification of topology.
We can also consider the change of the definition of $w$ to elucidate the difference of the parameter $w$ where the left and right edge modes appear. We discuss the possibility of different definitions of $w$ and the extension of the bulk-edge correspondence for such $w$ in the following section.
.9 Change of definition of the parameter $w$ according to the nonlinearity in more general systems.
@@
$i\partial_{t}\Psi_{j}(x)=\sum_{l,x^{\prime}}H_{jl}\left(x,x^{\prime};f(\Psi(x))\right)\Psi_{l}(x^{\prime}),$ (S86)
where $f(\Psi(x))$ is a nonlinear function of $\Psi_{j}(x)$ ($j=1,\cdots,M$ with $M$ being the internal degree of freedom), and $H$ is a matrix parametrized by $f(\Psi(x))$, it may be better to fix $f(\Psi(x))$ as $f(\Psi(x))=w$ instead of the amplitude.
We first show that fixing $f(\Psi(x))=w$ enables us to calculate the nonlinear eigenequation (Supplementary Eq. (S86)). The wavenumber-space description of the nonlinear eigenequation corresponding to Supplementary Eq. (S86) becomes
$E(k)\psi_{j}(k)=\sum_{l}H_{jl}\left(k;f(\psi(k))\right)\psi_{l}(k).$ (S87)
If we focus on special solutions with $f(\psi(k))=w$ being fixed independently of $k$, the right-hand side reads $\sum_{l}H_{jl}\left(k;w\right)\psi_{l}(k)$. Therefore, the nonlinear eigenvalue problem becomes equivalent to the linear eigenvalue problem parametrized by $w$ and $k$, and thus one can calculate the eigenvalues and eigenvectors by solving linear equations. We note that the multiple of the obtained eigenvector $\psi(k)$ is also an eigenvector because of the linearity. Thus, the final step of the calculation of the nonlinear eigenvector is to find the constant $c(k)$ that satisfies
$f(c(k)\psi(k))=w.$ (S88)
Such $c(k)$ can be absent for certain $\psi(k)$ and $w$, while one can guarantee the existence of $c(k)$ satisfying Supplementary Eq. (S88) if the nonlinear function has desirable properties. Specifically, if $f$ is a continuous function, one can always find $c(k)$ for $w$ in the range of $\max_{c}\min_{\psi}f(c\psi)\leq w\leq\min_{c}\max_{\psi}f(c\psi)$.
We can also discuss the bulk-edge correspondence of the winding number under the fixed $f(\Psi(x))=w$. In particular, if we focus on the generalized nonlinear SSH model,
$i\partial_{t}\Psi_{A}(x)$ $=$ $(a+f(\Psi_{A}(x),\Psi_{B}(x)))\Psi_{B}(x)+d\Psi_{B}(x-1),$ (S89)
$i\partial_{t}\Psi_{B}(x)$ $=$ $(a+f(\Psi_{A}(x),\Psi_{B}(x)))\Psi_{A}(x)+d\Psi_{A}(x+1),$ (S90)
@@
$\Psi_{A}(x+1)=-\frac{a+f(\Psi_{A}(x),0)}{d}\Psi_{A}(x).$ (S91)
When $a+f(\Psi_{A}(x),0)$ is smaller (larger) than $d$, $|\Psi_{A}(x+1)|$ becomes smaller (larger) than $|\Psi_{A}(x)|$. This implies that there is a fixed point at $\Psi^{\prime}_{A}(x)$ satisfying $a+f(\Psi^{\prime}_{A}(x),0)=d$, and it can be stable if $\nu(w)$ is one for at $w=f(\Psi^{\prime}_{A}(x)+\delta,0)$ and zero at $w=f(\Psi^{\prime}_{A}(x)-\delta,0)$ with $\delta$ being a sufficiently small constant. If $\Psi^{\prime}_{A}(x)$ is a stable fixed point, there are nonlinear edge modes with the initial amplitude $\Psi_{A}(1)$ satisfying $\nu(f(\Psi_{A}(1),0))=1$, which indicates the bulk-edge correspondence of the nonlinear winding number. However, Supplementary Equation (S91) can exhibit bifurcations, and $\Psi^{\prime}_{A}(x)$ can be an unstable fixed point after the bifurcations. Therefore, the bulk-edge correspondence is only valid for sufficiently weak nonlinearity as in the original nonlinear SSH model.

### v3

@@ Abstract
While we have used the localization indicator $P$ (Supplementary Eq. (S109)) in the above numerical calculations, there are other possible definitions of the localization indicators. Specifically, we can consider the following definitions that distinguish the localization and anti–localization of zero modes $\Psi_{j}(x)$ in a global sense: (1) $P_{1}=w_{\rm edge}/w_{\rm ave}$ with $w_{\rm ave}=\sum_{x,i}|\Psi_{i}(x)|^{2}/L$ being the averaged amplitude, (2) $P_{2}=w_{\rm edge}/w(x)$ with $w(x)=\sum_{i}|\Psi_{i}(x)|^{2}$ being the amplitude and $x$ denoting a bulk site that we determine in hand, (3) $P_{3}=w_{\rm edge}/w_{\rm max}$ with $w_{\rm max}=\max_{x}\left(\sum_{i}|\Psi_{i}(x)|^{2}\right)$ being the maximum amplitude. Unfortunately, these localization indicators can be inconsistent with the nonlinear winding number at $a<2$, where the bulk–edge correspondence still holds true in the nonlinear SSH model. Supplementary Figure S4 shows the localization indicators of the nonlinear eigenvectors with the minimum absolute values of nonlinear eigenvalues. At $a=1.5$, all the indicators correspond to the localization and anti–localization of zero modes predicted from the nonlinear winding number. In contrast, at $a=1.9$, the indicators do not correspond to the nonlinear winding number. In this parameter regime, the bulk–edge correspondence is confirmed in semi–infinite systems (Fig. 2 in the main text). Therefore, the localization indicators $P_{1}$, $P_{2}$, and $P_{3}$ are unsuitable to discuss the bulk–edge correspondence in finite nonlinear systems.
.11 Robustness of edge modes and chaos transitions against spatial disorders
We here confirm that topological edge modes and their chaos transitions are robust against disorders in the nonlinear SSH model. For simplicity, we consider the spatial inhomogeneity of $a$ and $d$ and fix the other parameters $b=c=-1$ independently of the location. We note that if we introduce disorders in $b$ and $c$, we can renormalize their effects into those of $a$ and $d$. We introduce the disorder in $a$, $d$ by setting $a(x)$, $d(x)$ at each location $x$ as $(1+\Delta a(x))a_{\rm ave}$, $(1+\Delta d(x))d_{\rm ave}$, where $\Delta a(x)$ and $\Delta d(x)$ are randomly determined from a uniform distribution $\left[-\delta,\delta\right]$. We fix the expectation value of $d(x)$ as $d_{\rm ave}=1$. We calculate the spatial dynamics similar to Eq. (7) in the main text at $\delta=0.01$ and $0.05$.
Figure S5: Bifurcation plots and Lyapunov exponents under the existence of disorder. a,b, Bifurcation plots under the existence of disorder. We fix the parameters $b=c=-1$ and $d_{\rm ave}=1$, and change $a_{\rm ave}$. We set the strength of disorder as $\delta=0.01$ in panel a and $\delta=0.05$ in panel b. The red lines at $a_{\rm ave}=1$ and $a_{\rm ave}=2$ represent the bifurcation points in the disorder-free case. The third red line shows the chaos transition point obtained from the numerical calculation of the Lyapunov exponents in panels c and d. We obtain almost convergent solutions at $a_{\rm ave}<2$, which indicates the robustness of topological edge modes against disorders. One can confirm the period–doubling bifurcation around $a_{\rm ave}\sim 2$ in panel a, while the other period–doubling bifurcations seen in the disorder–free case disappear. We note that at large $a_{\rm ave}$, the spatial dynamics diverge to infinity, and thus there are no data points. c,d, Lyapunov exponents under the existence of disorder. We use the same parameters as in panels a and b and set the strength of disorder as $\delta=0.01$ in panel c and $\delta=0.05$ in panel d. For both strengths of disorders, we obtain positive Lyapunov exponents (above the red lines) at any larger $a_{\rm ave}$ than critical values, which indicates the robustness of the chaos transition against disorder. At large $a_{\rm ave}$, since the spatial dynamics diverge to infinity, the Lyapunov exponents also diverge.
Supplementary Figure S5 shows the bifurcation diagram at each strength of the disorder. At $a_{\rm ave}<2$, the convergent behavior to $|\Psi_{A}(x)|\sim\sqrt{(a-b)/d}$ seems to remain if we permit the fluctuation of $\Psi_{A}(x)$ comparable to that of $a(x)$. We find that the period–doubling bifurcations disappear at large disorder $\delta$. At large $a_{\rm ave}$, the spatial dynamics becomes very noisy, which implies the spatial chaos of zero modes.
We also calculate the Lyapunov exponents under the existence of disorder (Supplementary Fig. S5c,d) as in Supplementary Fig. S1. We confirm that the Lyapunov exponent becomes positive at a certain $a_{\rm ave}$, and thus the spatial dynamics exhibits a chaos transition. Therefore, the chaos transition and the associated breakdown of the bulk–edge correspondence are also robust against disorders. We note that the periodic solutions after the chaos transition, i.e., the windows of chaos disappear under the existence of disorder. Such disorder–induced chaos is consistent with previous research Mayer1981 on the discrete nonlinear dynamics with noises.
In some nonlinear models, eigenvalues of topological edge modes can be shifted from zero even without disorders Tuloup2020 . In contrast, the present result indicates that the nonlinear eigenvalue of edge modes in the nonlinear SSH model is zero and immune to disorders and nonlinearity. This is because of the sublattice symmetry (Eq. (18) in Methods), which guarantees the symmetry of the band structure and the zero eigenvalue of edge modes (at least in the linear case).
.12 Stability of edge modes in nonlinear topological mechanics.
While we have discussed the destabilization of edge modes by chaos transitions in the main text, some nonlinear topological insulators do not exhibit such chaos transitions at arbitrary strength of the nonlinearity. We here analyze nonlinear topological mechanics composed of elliptic gears that are studied in Ref. Ma2023 , and show that the absence of chaos transitions is related to the linear stability in the spatial dynamics. The dynamics of that nonlinear topological mechanics is described as follows:
$i\partial_{t}\Psi_{A}(x)$ $=$ $f_{\epsilon}(\Psi_{B}(x))-f_{-\epsilon}(\Psi_{B}(x-1)),$ (S114)
$i\partial_{t}\Psi_{B}(x)$ $=$ $f_{\epsilon}(\Psi_{A}(x))-f_{-\epsilon}(\Psi_{A}(x+1)),$ (S115)
where $\Psi_{A,B}(x)$ are the complex field variable. $f_{\epsilon}(\Psi_{A,B}(x))$ is the nonlinear function defined as
$f_{\epsilon}(\Psi)$ $=$ $s_{\epsilon}({\rm Re}\,\Psi)+is_{\epsilon}({\rm Im}\,\Psi),$ (S116)
[…]
with $a$ being the semi major–axis of the elliptic gears and ${\epsilon}$ being their eccentricity.
Figure S6: Cobweb plots of the spatial dynamics of a topological mechanics. We present the cobweb plots of the spatial dynamics in Supplementary Eq. (S118). The orange (resp. blue) curves show $y=f_{\epsilon}(x)$ (resp. $y=f_{-{\epsilon}}(x)$). The red polylines represent the spatial dynamics. If we start from $\Psi_{A}(1)<\pi$, $\Psi_{A}(x)$ must be converged to zero at any $e$, which indicates the parameter independence of the stability of the fixed point at $\Psi_{A}(x)=0$. We fix the parameter $a=1$ and use ${\epsilon}=0.5$ in panel a and ${\epsilon}=0.9$ in panel b.
As in Eq. (7) in the main text, we focus on the zero nonlinear eigenvectors and derive the nonlinear dynamics describing their spatial distribution. Assuming $\Psi_{B}(x)=0$, one can determine $\Psi_{A}(x+1)$ from $\Psi_{A}(x)$ so that they satisfy
$f_{\epsilon}(\Psi_{A}(x))=f_{-{\epsilon}}(\Psi_{A}(x+1)).$ (S118)
Such spatial dynamics are captured by cobweb plots in Supplementary Fig. S6. Fortunately, the value of $\Psi_{A}(x+1)$ is uniquely determined at any parameters.
We investigate the spatial dynamics of zero modes at different ${\epsilon}$ and find that the fixed point at $\Psi_{A}(x)=0$ is always stable unless the bulk topology is changed. We confirm this from the fact that $\Psi_{A}(x+1)$ is smaller than $\Psi_{A}(x)$ for $\Psi_{A}(x)<\pi$ and ${\epsilon}>0$. Therefore, the zero mode must be decayed to zero in the thermodynamic limit $x\rightarrow\infty$, which indicates the absence of the chaos transition. The situation is similar in other nonlinear models in, e.g., Ref. Ma2023 , where the nonlinearity–induced topological modes are linearly stable at arbitrary parameters.
@@
.13 Linear stability analysis of the fixed points in the dynamical system describing zero modes of the extended nonlinear SSH model.
@@
Since the dynamical system in Eq. (9) in the main text can have three qualitatively different fixed points and limit cycles at $|\psi_{\rm fix}|^{2}=0$, $(d-a-\alpha)/b$, and $(\alpha-a)/b$, we below calculate the linear stability of these fixed points. We note that the nonzero fixed points, $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$ and $(\alpha-a)/b$, correspond to the topological edge modes that appear in the nonlinearity–induced topological phase because they converge to nonzero amplitude in the limit of $x\rightarrow\infty$. Below we focus on the parameter region $a,d,\alpha>0$, while we can conduct a similar analysis in the cases of different choices of the signs of the parameters.
.13.1 Stability analysis of the fixed point at $\psi_{\rm fix}=0$.
@@
Thus, one can determine the linear stability of the fixed point by comparing $a$ and $\alpha$. In the case of $a<\alpha$, the fixed point is linearly stable and has a two–dimensional stable manifold. In this parameter region, the winding number becomes $\nu_{\rm NL}=2$, which corresponds to the dimension of the stable manifold. In contrast, when $a$ is larger than $\alpha$, the fixed point is fully unstable. Then, the winding number becomes $\nu_{\rm NL}=0$, which is also equal to the dimension of the stable manifold.
@@
and thus confirm $|\lambda_{-}|>1$ and $|\lambda_{+}|<1$, which indicates that the stable manifold is one–dimension.. In contrast, if we consider the parameter region $d<a+\alpha$, we obtain $d^{2}-4\alpha a<(d-2\alpha)^{2}$. Then, the sign of $2\alpha-d$ is determined by the sign of $\alpha-a$ as shown from the inequality
@@
Therefore, the dimension of the stable manifold is two in the case of $\alpha>a$ and zero in the case of $\alpha<a$. These results are consistent with the bulk–edge correspondence between the winding number and the dimension of the stable manifold.
.13.2 Stability analysis of the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$.
If we consider the parameter region $d<a+\alpha$, $\alpha<a$, and $b<0$, we obtain the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$. This fixed point corresponds to the topological edge modes that emerge by the nonlinearity–induced topological phase transition at $w=(d-a-\alpha)/b$. In this case, the eigenvalues of the matrix in Supplementary Eq. (S119) become
@@
We can show that these eigenvalues are real. Since the nonlinear winding number becomes $\nu_{\rm NL}=1$ at $w>(d-a-\alpha)/b$, we expect that the dimension of the stable manifold of the fixed point at $|\psi_{\rm fix}|^{2}=(d-a-\alpha)/b$ is one. However, we can confirm the breakdown of such bulk–edge correspondence by the period–doubling bifurcation as in the original nonlinear SSH model in Eqs. (4) and (5) in the main text.
@@
Therefore, $|\lambda_{+}|$ becomes smaller than one and thus the nonlinear winding number corresponds to the dimension of the stable manifold only in the case of $(a+\alpha)/d>2$. At the critical parameter $(a+\alpha)/d=2$, the dynamical system exhibits a period–doubling bifurcation. Such bifurcation induces the breakdown of the bulk–edge correspondence as in the original nonlinear SSH model in Eqs. (4) and (5) in the main text.
.13.3 Stability analysis of the fixed point at $|\psi_{\rm fix}|^{2}=(\alpha-a)/b$.
If we consider the case of $2\alpha>d$ and $\alpha<a$, we obtain a limit cycle described as $\psi(x)=\sqrt{(\alpha-a)/b}\exp(i(\Omega x+\theta))$ whose frequency $\Omega$ satisfies $\Omega=-2\alpha\cos\theta$. This limit cycle corresponds to the topological edge mode that emerges by the nonlinearity–induced topological phase transition at $w=(\alpha-a)/b$, and the nonlinear winding number is changed into $\nu_{\rm NL}=2$ at the transition point. The eigenvalues of the linearized matrix in Supplementary Eq. (S119) become
@@
Thus, we obtain $|\lambda_{+}|<1$ only at $a-d/2<2\alpha$. If we consider the case of $a+d/2>2\alpha$, the dimension of the stable manifold is less than two, which indicates the breakdown of the bulk–edge correspondence. This breakdown is also induced by the bifurcation of the fixed point. Therefore, the breakdown of the bulk–edge correspondence by the chaos transition is independent of the hopping range in nonlinear models.
.14 Increased number of anti–localized modes in the long–range nonlinear SSH model.
We here show that long–range hopping can also increase the number of anti–localized zero modes as well as topological localized modes. As discussed in the main text, we can regard the dimension of the stable manifold as the effective number of nonlinear zero eigenvectors. In the long–range nonlinear SSH model (Eq. (9) in the main text), nonzero fixed points can have a two–dimensional stable manifold, which implies the existence of two independent anti–localized zero modes (this also indicates the existence of two localized zero modes).
Figure S7: Two anti–localized zero modes in the long–range nonlinear SSH model. a Spatial distribution of two anti–localized zero modes are presented by the red and blue polylines. By comparing the absolute value of the zero modes at the edge $x=1$ with that in the bulk $x\rightarrow\infty$, we judge their anti–localization. These anti–localized zero modes have different values at the edge, and thus we distinguish them. We use the parameters $a=0.11$, $b=c=1$, $d=-1$, and $\alpha=0.8$. b The flow of the nonlinear dynamical system of the long–range nonlinear SSH model (Eq. (9) in the main text) and the orbits corresponding to anti–localized zero modes. The blue curved arrows show the flow of the nonlinear dynamical system in the spatial direction. The green square and red disks are the fixed points, where the colors represent the dimensions of their stable manifolds: one (two) dimensions for the green circle (red squares). One can estimate that the red squares have two–dimensional stable manifolds from the flow around them; starting from any neighbor point, the orbit converges to the fixed point. On the other hand, the perturbation around the green fixed point diverges except for one certain direction presented by the orange line, and thus the fixed point at origin has a one–dimensional stable manifold. The blue and red curves show the orbits corresponding to the anti–localized zero modes in panel a. Since neither of the curves is a part of the other, they correspond to independent localized modes and indicate that the values of anti–localized modes at the edge $\Psi(1)$ have two degrees of freedom. We use the same parameters as in panel a.
We numerically demonstrate the existence of two anti–localized zero modes in the long–range nonlinear SSH model. We consider the parameters $a=0.11$, $b=c=1$, $d=-1$, and $\alpha=0.8$. In this case, the nonlinear winding number becomes $\nu(w)=1$ at the amplitude $w<0.09$, and $\nu(w)=2$ at $w>0.09$. We calculate the spatial distribution of zero modes by using the nonlinear dynamical system in Eq. (9) in the main text, starting from two different initial conditions $(\Psi_{A}(1),\Psi_{A}(2))=(0.1,\pm 0.1)$. From both of the initial conditions, $\Psi_{A}(x)$ converges to $\Psi_{A}(x)\rightarrow 0.3$, which is larger than the initial amplitude $0.1$ (cf. Supplementary Fig. S7a). Therefore, we obtain two anti–localized zero modes. This increase in the number of anti–localized zero modes corresponds to the two–dimensional stable manifold of the fixed point at $(\Psi_{A}(x),\Psi_{A}(x+1))=(0.3,0.3)$ in Supplementary Fig. S7b (two dimensions of the stable manifold are analytically shown in Supplementary Note 7).
.15 Correspondence between the nonlinear winding number and the eigenvalues of a state-dependent transfer matrix.
We have also considered more general nonlinear systems whose zero modes are described by Eq. (11) in the main text. We here show that the eigenvalues of $T(g(\bm{\psi}_{1}(x)))$ in Eqs. (11) and (12) in the main text are related to the nonlinear winding number, which is the origin of the bulk–edge correspondence. Specifically, if we consider a fixed value of $g(\bm{\psi}(k))=w$ independently of the wavenumber $k$, the nonlinear winding number
@@
.16 Possible $\mathbb{Z}\times\mathbb{Z}$ classification in the nonlinear SSH model.
If we consider the case of $b\neq c$ in the nonlinear SSH model (Eqs. (4) and (5) in the main text), edge modes localized at the left and right boundaries can appear at different critical amplitudes. To show that, we consider both left and right semi–infinite systems, $x\leq-1$ and $x\geq 1$, respectively. While the spatial distribution of zero modes in the right semi–infinite system is described by
@@
(the same as Eq. (7) in the main text), that in the left semi–infinite system is described by
@@
and $\Psi_{A}(x)=0$. The difference between $b$ and $c$ leads to the different transition points in Supplementary Eqs. (S138) and (S139); the edge mode in the right semi–infinite system appears at $w=(d-a)/b$, while that in the left semi–infinite system appears at $w=(d-a)/c$. Therefore, the left edge mode and right edge mode have different critical amplitudes.
@@
whose transition points are consistent with the amplitude where the left and right edge modes appear. Such a $\mathbb{Z}\times\mathbb{Z}$ classification may be analogous to that in non-Hermitian systems, where the point-gap topology of one–dimensional systems is also classified by $\mathbb{Z}\times\mathbb{Z}$ topological invariants under the existence of the sublattice symmetry Kawabata2019 . However, the nonlinear SSH model considered here is a conservative system and preserves the energy in its time evolution. Therefore, nonlinearity in itself can induce non-Hermitian-like effects in the classification of topology.
We can also consider the change of the definition of $w$ to elucidate the difference of the parameter $w$ where the left and right edge modes appear. We discuss the possibility of different definitions of $w$ and the extension of the bulk–edge correspondence for such $w$ in the following section.
.17 Change of definition of the parameter $w$ according to the nonlinearity in more general systems.
@@
$i\partial_{t}\Psi_{j}(x)=\sum_{l,x^{\prime}}H_{jl}\left(x,x^{\prime};g(\Psi(x))\right)\Psi_{l}(x^{\prime}),$ (S143)
where $g(\Psi(x))$ is a nonlinear function of $\Psi_{j}(x)$ ($j=1,\cdots,M$ with $M$ being the internal degree of freedom), and $H$ is a matrix parametrized by $g(\Psi(x))$, it may be better to fix $g(\Psi(x))$ as $g(\Psi(x))=w$ instead of the amplitude.
We first show that fixing $g(\Psi(x))=w$ enables us to calculate the nonlinear eigenequation (Supplementary Eq. (S143)). The wavenumber–space description of the nonlinear eigenequation corresponding to Supplementary Eq. (S143) becomes
$E(k)\psi_{j}(k)=\sum_{l}H_{jl}\left(k;g(\psi(k))\right)\psi_{l}(k).$ (S144)
If we focus on special solutions with $g(\psi(k))=w$ being fixed independently of $k$, the right-hand side reads $\sum_{l}H_{jl}\left(k;w\right)\psi_{l}(k)$. Therefore, the nonlinear eigenvalue problem becomes equivalent to the linear eigenvalue problem parametrized by $w$ and $k$, and thus one can calculate the eigenvalues and eigenvectors by solving linear equations. We note that the multiple of the obtained eigenvector $\psi(k)$ is also an eigenvector because of the linearity. Thus, the final step of the calculation of the nonlinear eigenvector is to find the constant $c(k)$ that satisfies
$g(c(k)\psi(k))=w.$ (S145)
Such $c(k)$ can be absent for certain $\psi(k)$ and $w$, while one can guarantee the existence of $c(k)$ satisfying Supplementary Eq. (S145) if the nonlinear function has desirable properties. Specifically, if $g$ is a continuous function, one can always find $c(k)$ for $w$ in the range of $\max_{c}\min_{\psi}g(c\psi)\leq w\leq\min_{c}\max_{\psi}g(c\psi)$.
We can also discuss the bulk–edge correspondence of the winding number under the fixed $f(\Psi(x))=w$. In particular, if we focus on the generalized nonlinear SSH model,
$i\partial_{t}\Psi_{A}(x)$ $=$ $(a+g(\Psi_{A}(x),\Psi_{B}(x)))\Psi_{B}(x)+d\Psi_{B}(x-1),$ (S146)
$i\partial_{t}\Psi_{B}(x)$ $=$ $(a+g(\Psi_{A}(x),\Psi_{B}(x)))\Psi_{A}(x)+d\Psi_{A}(x+1),$ (S147)
@@
$\Psi_{A}(x+1)=-\frac{a+g(\Psi_{A}(x),0)}{d}\Psi_{A}(x).$ (S148)
When $a+g(\Psi_{A}(x),0)$ is smaller (larger) than $d$, $|\Psi_{A}(x+1)|$ becomes smaller (larger) than $|\Psi_{A}(x)|$. This implies that there is a fixed point at $\Psi^{\prime}_{A}(x)$ satisfying $a+g(\Psi^{\prime}_{A}(x),0)=d$, and it can be stable if $\nu(w)$ is one for at $w=g(\Psi^{\prime}_{A}(x)+\delta,0)$ and zero at $w=g(\Psi^{\prime}_{A}(x)-\delta,0)$ with $\delta$ being a sufficiently small constant. If $\Psi^{\prime}_{A}(x)$ is a stable fixed point, there are nonlinear edge modes with the initial amplitude $\Psi_{A}(1)$ satisfying $\nu(g(\Psi_{A}(1),0))=1$, which indicates the bulk–edge correspondence of the nonlinear winding number. However, Supplementary Equation (S148) can exhibit bifurcations, and $\Psi^{\prime}_{A}(x)$ can be an unstable fixed point after the bifurcations. Therefore, the bulk–edge correspondence is only valid for sufficiently weak nonlinearity as in the original nonlinear SSH model.